# Upgrade of Super-Kamiokande with Gadolinium

### v1

@@ 5.2 Purification Method
To meet the strict radio-purity requirements for SK-Gd, a multi-step purification method was developed to produce ultra-pure $\rm Gd_{2}(\rm SO_{4})_{3}\cdot\rm 8H_{2}O$ starting from gadolinium oxide ($\rm Gd_{2}\rm O_{3}$). The raw material $\rm Gd_{2}\rm O_{3}$, with >>99.99% purity (Gd2O3/TREO), was selected with special care to minimize thorium and uranium concentrations.
@@ 7.1 Search for DSNB
The properties of the prompt events, such as vertex, direction, and energy, are reconstructed using the same algorithms as in the SK solar neutrino analysis. The method of selecting events for DSNB to reduce background has been described in detail in several papers such as [21] . These reductions have reduced the remaining background to a level approximately equivalent to the expected DSNB signal. Figure 9 shows the reconstructed energy distribution of the prompt events after all cuts are applied. We concluded that there was no significant DSNB signal, although a slight excess was seen in some bins.
@@ 7.2 Neutrons from Atmospheric Neutrinos
When gamma rays and a single neutron are observed resulting from this interaction, it becomes a serious background in the DSNB search because the reconstructed gamma ray energy is similar to the positron energy in IBD from DSNB, and it is difficult to distinguish them from the DSNB signal. Therefore, research was also conducted focusing on the NCQE event itself. In this study, several simulations were compared with data regarding the PMT hit pattern and reconstructed energy of the prompt gamma ray signal, and the neutron multiplicity obtained from the delayed signal. The major difference observed here was the hit pattern from the prompt gamma rays, and the result showed that the data for prompt gamma ray multiplicity was significantly less than in the simulation previously used. Further investigation revealed that the cause of this difference was nuclear de-excitation, which could be improved by using the Precompound model implemented in Geant4. This is consistent with the results of an external (non-SK) experiment in which a water target was irradiated with a neutron beam and the generated gamma rays were measured with a germanium detector.

### v2

@@ 5.2 Purification Method
To meet the strict radio-purity requirements for SK-Gd, a multi-step purification method was developed to produce ultra-pure $\rm Gd_{2}(\rm SO_{4})_{3}\cdot\rm 8H_{2}O$ starting from gadolinium oxide ($\rm Gd_{2}\rm O_{3}$). The raw material $\rm Gd_{2}\rm O_{3}$, with >>99.99% purity (Gd2O3/Total Rare Earth Oxide), was selected with special care to minimize thorium and uranium concentrations.
@@ 7.1 Search for DSNB
The properties of the prompt events, such as vertex, direction, and energy, are reconstructed using the same algorithms as in the SK solar neutrino analysis [35]. The method of selecting events for DSNB to reduce background has been described in detail in [12, 13, 33]. These reductions have reduced the remaining background to a level approximately equivalent to the expected DSNB signal. Figure 9 shows the reconstructed energy distribution of the prompt events after all cuts are applied. We concluded that there was no significant DSNB signal, although a slight excess was seen in some bins.
@@ 7.2 Neutrons from Atmospheric Neutrinos
When gamma rays and a single neutron are observed resulting from this interaction, it becomes a serious background in the DSNB search because the reconstructed gamma ray energy is similar to the positron energy in IBD from DSNB, and it is difficult to distinguish this background from the DSNB signal. Therefore, the NCQE signal itself was investigated. In this study, several simulations were compared with data regarding the PMT hit pattern and reconstructed energy of the prompt gamma ray signal, and the neutron multiplicity obtained from the delayed signal. The major difference observed here was the hit pattern from the prompt gamma rays, and the result showed that the data for prompt gamma ray multiplicity was significantly less than in the simulation previously used. Further investigation revealed that the cause of this difference was nuclear de-excitation, which could be improved by using the Precompound model to simulate the emission of baryons from recently struck, highly excited nuclei, as implemented in Geant4 [39]. This is consistent with the results of an external (non-SK) experiment in which a water target was irradiated with a neutron beam and the generated gamma rays were measured with a germanium detector.